OpenStudy (anonymous):

The half-life of a certain first order reaction is 15 minutes. What fraction of the original reactant concentration will remain after 2.0 hours??

OpenStudy (anonymous):

I know the half life formula for a first order reaction is \[t _{1/2} = \frac{ \ln(2) }{ }\]

OpenStudy (anonymous):

divided by k

OpenStudy (aaronq):

for a first order reaction, use this equation to find the decay constant: \[t _{1/2}=\frac{ \ln2 }{ k }\] then use it in the exponent growth/decay equation (which I'm sure you know)

OpenStudy (aaronq):

haha ok good

OpenStudy (anonymous):

the thing is we have not worked with the exponent growth/decay in this class.

OpenStudy (anonymous):

we've only had the integrated rate laws, half like formulas and arrhenius equation

OpenStudy (aaronq):

ohh okay, do you know what the equation is? do you know how to solve logarithms?

OpenStudy (aaronq):

okay, so since you've used the arrheius equation I'm guessing you know how to solve logarithms

OpenStudy (anonymous):

yes.

OpenStudy (aaronq):

so the equation is this: \[A=A _{0}e ^{-kt}\] A is the amount after whatever time elapsed Ao is the initial amount k is the constant t is the time elapsed

OpenStudy (aaronq):

so it's basically a plug and chug after that, but you have to take into account the differences in the time units

OpenStudy (anonymous):

But I dont have the initial concentration...

OpenStudy (aaronq):

you can assume the initial to be 1 or 100

OpenStudy (aaronq):

because it's asking you for the difference (fraction)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

let me do it and Ill post what I get

OpenStudy (aaronq):

k

OpenStudy (anonymous):

alright, Im lost. I thoufht I had it. this is what I did \[\ln \frac{ x }{ 100 }=-0.046(120)\]

OpenStudy (aaronq):

the answer should be 1/256 because 15 min is your half-life, then in 2 hours it has decayed 8 times (1/2)^8 = 0.00390625 = 1/256

OpenStudy (aaronq):

if you want il work it out the formal way

OpenStudy (anonymous):

yeah, the answer is 1/256 but i didn't know how to get there without the initial concentration

OpenStudy (aaronq):

okay.. so k=ln2/(1/4)=4ln2=2.7725 A=100e^(-2.7725(2)) A=100e^(-5.545177444) A=100(3.90625x10^-3)= 0.390625 100/0.390625 = 256 1/256

OpenStudy (aaronq):

i took 15 mins as 1/4 of an hour just to simplify things

OpenStudy (anonymous):

ok thanks! cuz I was jsut about to ask you where you got the 1/4 from

OpenStudy (aaronq):

no problem!